OpenStudy (anonymous):

Bart opened a credit card at a department store with an APR of 19.99% compounded monthly. What is the APY on this credit card?

OpenStudy (anonymous):

So formula is APY=(1+P)^n-1 and I did

OpenStudy (anonymous):

APY=(1+0.1999/12)^12-12

OpenStudy (anonymous):

and got wait I did 12 -1 instead of 12-12

OpenStudy (alexwee123):

well first of all do you know what APY stands for?

OpenStudy (anonymous):

Annual Percentage Yield Ok so I got 1.01665833333333 what did I do wrong?

OpenStudy (anonymous):

because that is not one of my answer choices

OpenStudy (alexwee123):

what are the choices? :/

OpenStudy (anonymous):

39.98% 26.81% 21.93% 1.67%

OpenStudy (anonymous):

I am supposed to solve with APY=(1+0.1999/12)^12-1

OpenStudy (alexwee123):